Some Notes on Alternating Power Sums of Arithmetic Progressions

András Bazsó Institute of Mathematics University of Debrecen

and

MTA-DE Research Group “Equations Functions and Curves” Hungarian Academy of Sciences and University of Debrecen

P. O. Box 400 H-4002 Debrecen

Hungary

bazsoa@science.unideb.hu

István Mező Department of Mathematics

Nanjing University of Information Science and Technology

No. 219 Ningliu Rd. Pukou, Nanjing, Jiangsu

PR China

istvanmezo81@gmail.com

Abstract

We show that the alternating power sum

r^{n} − (m + r)^{n} + (2m + r)^{n} − ⋯ + (−1)^{ℓ−1} ((ℓ − 1) m + r)^{n}

can be expressed in terms of Stirling numbers of the first kind and r-Whitney numbers of the second kind. We also prove a necessary and sufficient condition for the integrality of the coefficients of the polynomial extensions of the above alternating power sum.

1 Introduction

Power sums and alternating power sums of consecutive numbers are widely investigated objects in the literature of combinatorics and number theory. It is well known, among others, that the sum of the n-th power of the first ℓ − 1 positive integers

S_{n}(ℓ) := 1^{n} + 2^{n} + ⋯ + (ℓ − 1)^{n}

is closely connected to the Bernoulli polynomials B_{n}(x) via the identity

S_{n}(ℓ) = \frac{1}{n + 1} (B_{n+1}(ℓ) − B_{n+1}) ,

where the polynomials B_{n}(x) are defined by the generating series

\frac{te^{tx}}{e^{t} − 1} = ∑_{k=0}^{∞} B_{k}(x)\frac{t^{k}}{k!}

and B_{n+1} = B_{n+1}(0).

It is also well known that the alternating power sum

T_{n} (ℓ) := −1^{n} + 2^{n} − ⋯ + (−1)^{ℓ−1}(ℓ − 1)^{n}

can be expressed by means of the classical Euler polynomials E_{n}(x) via:

T_{n} (ℓ) = \frac{E_{n}(0) + (−1)^{ℓ−1}E_{n}(ℓ)}{2},

where the classical Euler polynomials E_{n}(x) are usually defined by the generating function

\frac{2e^{xt}}{e^{t} + 1} = ∑_{k=0}^{∞} E_{k}(x)\frac{t^{k}}{k!} (|t| < π).

For the properties of Bernoulli and Euler polynomials which will be often used in this paper, sometimes without special reference, we refer to the paper of Brillhart [9] and the book of Abramowitz and Stegun [1].

Let ℓ > 1, m ≠ 0, r be integers with gcd(m, r) = 1 and consider the following sums

S_{m,r}^{n} (ℓ) = r^{n} + (m + r)^{n} + (2m + r)^{n} + ⋯ + ((ℓ − 1) m + r)^{n} ,
T_{m,r}^{n} (ℓ) := r^{n} − (m + r)^{n} + (2m + r)^{n} − ⋯ + (−1)^{ℓ−1} ((ℓ − 1) m + r)^{n} .

Bazsó et al. [5] showed that S_{m,r}^{n} (ℓ) can be extended to the following polynomial in x:

S_{m,r}^{n} (x) = \frac{m^{n}}{n + 1} (B_{n+1} (x + \frac{r}{m}) − B_{n+1} (\frac{r}{m})) .  (1)

Using a different approach, Howard [11] also obtained relation (1) together with its analogue for T_{m,r}^{n} (ℓ):

T_{m,r}^{n} (ℓ) = \frac{m^{n}}{2} (E_{n} (\frac{r}{m}) + (−1)^{ℓ−1}E_{n} (ℓ + \frac{r}{m})) ,  (2)

whence, the following polynomial extensions arise for T_{m,r}^{n} (ℓ):

T_{m,r}^{n+}(x) = \frac{m^{n}}{2} (E_{n} (\frac{r}{m}) + E_{n} (x + \frac{r}{m})) ,  (3)

T_{m,r}^{n−}(x) = \frac{m^{n}}{2} (E_{n} (\frac{r}{m}) − E_{n} (x + \frac{r}{m})) .  (4)

Clearly, for positive integer values x, we have T_{m,r}^{n+}(x) = T_{m,r}^{n} (x) if x is odd, and T_{m,r}^{n−}(x) = T_{m,r}^{n} (x) if x is even.

For related diophantine results on the polynomials S_{m,r}^{n} (x) , T_{m,r}^{n+}(x), and T_{m,r}^{n−}(x) see [3, 6, 7, 8, 10, 12, 15] and the references given there. For results on the decomposition of these polynomials we refer to the papers [2, 5, 8].

In a recent paper [4], the present authors investigated the coefficients of S_{m,r}^{n} (x). We showed that these coefficients can be given in terms of the Stirling numbers of the first kind and r-Whitney numbers of the second kind. Moreover, we proved that S_{m,r}^{n} (x) ∈ ℤ[x] if and only if m is divisible by F (n), where F (n) is the sequence with first few terms

2, 6, 2, 30, 6, 42, 6, 30, 10, 66, 6, 2730, …

(cf. \underline{A144845} in Sloane’s OEIS [14]). We [4] also gave an implicit formula for F (n).

The aim of this note is to give analogues of our results [4] on S_{m,r}^{n} (x) for the alternating case.

2 An explicit formula for the alternating sum T_{m,r}^{n}(ℓ)

From (2) we know that the alternating sum T_{m,r}^{n}(ℓ) can be expressed in terms of ℓ and the Euler polynomials. We give an explicit formula for T_{m,r}^{n}(ℓ) without the Euler polynomials included. To do this we need the following lemma.

Lemma 1. For all ℓ ≥ 1 and k ≥ 0 we have that

∑_{x=0}^{ℓ−1}(−1)^{x}x\underline{k} = k!\frac{(−1)^{k}}{2^{k+1}} (1 + (−1)^{ℓ+1} ∑_{i=0}^{k} \binom{ℓ}{i}(−2)^{i}) .

Here x\underline{k} = x(x − 1) ⋯ (x − k + 1) is the falling factorial.

Proof. The idea we use here is due to Felix Marin. We found out about his idea on the Mathematics Stack Exchange forum [17].

First note that

∑_{x=0}^{ℓ−1}(−1)^{x}x\underline{k} = k! ∑_{x=0}^{ℓ−1}(−1)^{x}\binom{x}{k}.  (5)

Then we use the integral representation

\binom{x}{k} = ∮_{|z|<1} \frac{(1 + z)^{x}}{z^{k+1}} \frac{dz}{2πi}.

Substituting this into (5) the summation can already be done. We have the intermediate result that

∑_{x=0}^{ℓ−1}(−1)^{x}x\underline{k} = k! ∮_{|z|<1} \frac{1}{z^{k+1}} \frac{(−1)^{ℓ+1}(1 + z)^{ℓ} + 1}{2 + z} \frac{dz}{2πi}.  (6)

The path integral on the right can be calculated by Cauchy’s residue theorem:

∮_{|z|<1} \frac{1}{z^{k+1}} \frac{(−1)^{ℓ+1}(1 + z)^{ℓ} + 1}{2 + z} \frac{dz}{2πi} =

Res_{z=0} \frac{1}{(2 + z)z^{k+1}} + (−1)^{ℓ+1}Res_{z=0} \frac{(1 + z)^{ℓ}}{(2 + z)z^{k+1}} .

The first residue is easy to determine:

Res_{z=0} \frac{1}{(2 + z)z^{k+1}} = \frac{(−1)^{k}}{2^{k+1}} (k ≥ 0).  (7)

The calculation of the second residue can be traced back to the first one by expanding (1+z)^{ℓ} by the binomial theorem:

Res_{z=0} \frac{(1 + z)^{ℓ}}{(2 + z)z^{k+1}} = ∑_{i=0}^{ℓ} \binom{ℓ}{i}Res_{z=0} \frac{z^{i}}{(2 + z)z^{k+1}} .

Note that if i ≥ k + 1 the function becomes analytic at z = 0 and the residue disappears. Hence, recalling (7),

Res_{z=0} \frac{(1 + z)^{ℓ}}{(2 + z)z^{k+1}} = ∑_{i=0}^{k} \binom{ℓ}{i}\frac{(−1)^{k−i}}{2^{k+1−i}} = \frac{(−1)^{k}}{2^{k+1}} ∑_{i=0}^{k} \binom{ℓ}{i}(−2)^{i}.

This and (7) together gives the result.

We recall the definition of the r-Whitney numbers W_{m,r}(n, k) given by Mező [13]:

(mx + r)^{n} = ∑_{k=0}^{n} m^{k}W_{m,r}(n, k)x\underline{k}.  (8)

Theorem 2. If ℓ ≥ 1 then the sum T_{m,r}^{n}(ℓ) can be expressed as follows:

T_{m,r}^{n}(ℓ) = (1 + (−1)^{ℓ+1}) ∑_{k=0}^{n} C_{m,r,n,k}+

∑_{j=1}^{n} ℓ^{j} ((−1)^{ℓ+1} ∑_{k=0}^{n} C_{m,r,n,k} ∑_{i=j}^{k} \frac{(−2)^{i}}{i!} S_{1}(i, j)) .

Here

C_{m,r,n,k} = k!\frac{(−1)^{k}}{2^{k+1}} m^{k}W_{m,r}(n, k),

and W_{m,r}(n, k) is an r-Whitney number.

Proof. We can see that it is enough to multiply both sides of (8) by (−1)^{x} and sum from x = 0, 1, …, ℓ − 1 to get back T_{m,r}^{n}(ℓ). Hence

T_{m,r}^{n}(ℓ) = ∑_{k=0}^{n} m^{k}W_{m,r}(n, k) ∑_{x=0}^{ℓ−1}(−1)^{x}x\underline{k}.

By the previous lemma we now have that

T_{m,r}^{n}(ℓ) = ∑_{k=0}^{n} m^{k}W_{m,r}(n, k)k!\frac{(−1)^{k}}{2^{k+1}} (1 + (−1)^{ℓ+1} ∑_{i=0}^{k} \binom{ℓ}{i}(−2)^{i}) .  (9)

Our original goal is to find the coefficients of ℓ in this expression. It is immediate that the constant term equals to (when i = 0)

C := (1 + (−1)^{ℓ+1}) ∑_{k=0}^{n} m^{k}W_{m,r}(n, k)k!\frac{(−1)^{k}}{2^{k+1}} .  (10)

The other coefficients of ℓ can be determined by expanding the \binom{ℓ}{i} binomial coefficients with the aid of the Stirling numbers of the first kind:

\binom{ℓ}{i} = \frac{1}{i!} ∑_{j=0}^{i} S_{1}(i, j)ℓ^{j}.

Here, the index of the sum runs to n (this is the maximal value j can ever attain), because S_{1}(i, j) = 0 if j > i. So we can factor out ℓ^{j} in (9):

T_{m,r}^{n}(ℓ) = C + ∑_{j=0}^{n} ℓ^{j} ∑_{k=0}^{n} m^{k}W_{m,r}(n, k)k!\frac{(−1)^{k}}{2^{k+1}} (−1)^{ℓ+1} ∑_{i=0}^{k} \frac{(−2)^{i}}{i!} S_{1}(i, j).

Noting that i runs from j, and recalling the definition of the constants C_{m,r,n,k} we finish the proof.

3 The integrality of the coefficients of the polynomials T_{m,r}^{n+}(x) and T_{m,r}^{n−}(x)

In this section, let m, r, n be integers with m ≠ 0, r coprime and n > 0.

Theorem 3. For all m, r and n, both 2(n + 1)T_{m,r}^{n+}(x) and 2(n + 1)T_{m,r}^{n−}(x) are in ℤ[x].

Our Theorem 3 follows from the following result.

Lemma 4. For all m and n we have (n + 1)m^{n}E_{n} (\frac{x}{m}) ∈ ℤ[x].

Proof. This is part of a result of Sun [16, Lemma 2.2].

Proof of Theorem 3. By (3) we have

T_{m,r}^{n+}(x) = \frac{m^{n}}{2} (E_{n} (\frac{r}{m}) + E_{n} (x + \frac{r}{m})) =

= \frac{1}{2(n + 1)} [(n + 1)m^{n} (E_{n} (\frac{r}{m}) + E_{n} (\frac{mx + r}{m}))] . (11)

The expression in square brackets is a sum of two polynomials with integer coefficients by Lemma 4, whence 2(n + 1)T_{m,r}^{n+}(x) ∈ ℤ[x]. For 2(n + 1)T_{m,r}^{n−}(x), the proof is essentialy the same.

By the denominator of a polynomial P (x) ∈ ℚ[x] we mean the smallest positive integer d such that dP (x) ∈ ℤ[x]. An immediate consequence of Theorem 3 is that the denominators of T_{m,r}^{n+}(x) and T_{m,r}^{n−}(x) respectively, are divisors of 2(n + 1). In the sequel, we give a more precise description of these denominators.

Remark 5. It is well known (see, e.g., the paper of Brillhart [9] p. 46) that an Euler polynomial of even index has only integer coefficients, and that the denominator of an odd index Euler polynomial is a power of 2. By Lemma 4 with choice m = 1, the denominator of the n-th Euler polynomial is either 1 or a power of 2 which divides n + 1.

Now we state the main result of this section.

Theorem 6. All coefficients of the polynomials T_{m,r}^{n+}(x) and T_{m,r}^{n−}(x) are integers if and only if m is even.

Proof. First we consider the integrality of the coefficients of the polynomial T_{m,r}^{n+}(x). By (3), we observe that all these coefficients are integers if and only if m^{n} is divisible by the denominator of T_{m,r}^{n+}(x). Let this denominator be denoted by D. Clearly, D is the product of 2 and the denominator of E_{n}(x).

If n is even, then D = 2 by the above remark, and thus for even m all the coefficients of T_{m,r}^{n+}(x) are integers.

For odd n, by the same remark, the denominator of E_{n}(x) is a power of 2 which divides n + 1, say 2^{q}. Thus we have D = 2^{q+1}. Since 2^{q} ≤ n + 1 < 2^{n} for n > 1, it follows that n ≥ q + 1, thus D divides 2^{n} for n > 1. For n = 1, we have

T_{m,r}^{1+}(x) = \frac{mx}{2} − \frac{m}{2} + r.

Hence for even m, we have T_{m,r}^{n+}(x) ∈ ℤ[x].

The equivalence of the integrality of the coefficients of T_{m,r}^{n−}(x) and that m is even follows from a similar argument and from (4). This completes the proof.

4 Acknowledgments

The authors are grateful to the referee for her/his useful comments and suggestions.

The first author was supported by the Hungarian Academy of Sciences and by the OTKA grant NK104208. The second author was supported by the Scientific Research Foundation of Nanjing University of Information Science & Technology, by the project S8113062001 of the Startup Foundation for Introducing Talent of NUIST, and by the grant 11501299 of the National Natural Science Foundation for China.

References

[1] M. Abramowitz and I. A. Stegun, Handbook of Mathematical Functions, National Bureau of Standards, 1964.

[2] A. Bazsó, On alternating power sums of arithmetic progressions, Integral Transforms Spec. Funct. 24 (2013), 945–949.

[3] A. Bazsó, Polynomial values of (alternating) power sums, Acta Math. Hungar. 146 (2015), 202–219.

[4] A. Bazsó and I. Mező, On the coefficients of power sums of arithmetic progressions, J. Number Theory 153 (2015) 117–123.

[5] A. Bazsó, A. Pintér, and H. M. Srivastava, On a refinement of Faulhaber’s theorem concerning sums of powers of natural numbers, Appl. Math. Lett. 25 (2012), 486–489.

[6] A. Bazsó, D. Kreso, F. Luca, and Á. Pintér, On equal values of power sums of arithmetic progressions, Glas. Mat. Ser. III 47 (2012), 253–263.

[7] M. A. Bennett, A superelliptic equation involving alternating sums of powers, Publ. Math. Debrecen 79 (2011), 317–324.

[8] Y. F. Bilu, B. Brindza, P. Kirschenhofer, Á. Pintér, and R. F. Tichy, Diophantine equations and Bernoulli polynomials (with an Appendix by A. Schinzel), Compositio Math. 131 (2002), 173–188.

[9] J. Brillhart, On the Euler and Bernoulli polynomials, J. Reine Angew. Math. 234 (1969), 45–64.

[10] K. Győry and Á. Pintér, On the equation 1^{k} + 2^{k} + ⋯ + x^{k} = y^{n}, Publ. Math. Debrecen 62 (2003), 403–414.

[11] F. T. Howard, Sums of powers of integers via generating functions, Fibonacci Quart. 34 (1996), 244–256.

[12] D. Kreso and Cs. Rakaczki, Diophantine equations with Euler polynomials, Acta Arith. 161 (2013), 267–281.

[13] I. Mező, A new formula for the Bernoulli polynomials, Result. Math. 58 (2010), 329–335.

[14] N. J. A. Sloane, The On-Line Encyclopedia of Integer Sequences (OEIS), https://

    oeis.org.

[15] Cs. Rakaczki, On the Diophantine equation S_{m}(x) = g(y), Publ. Math. Debrecen 65 (2004), 439–460.

[16] Z.-W. Sun, Explicit congruences for Euler polynomials, in Number Theory, Dev. Math., Vol. 15, Springer, 2006, pp. 205–218.

[17] H. Yuen, Truncated alternating binomial sum, Question posted on Mathe-

    matics Stack Exchange, http://www.math.stackexchange.com/questions/887960/
    truncated-alternating-binomial-sum.

2010 Mathematics Subject Classification: Primary 11B68. Secondary 11B25, 11B73. Keywords: arithmetic progression, power sum, Euler polynomial, Stirling number, r-Whitney number.

(Concerned with sequence \underline{A144845}.)

Received April 25 2017; revised versions received May 1 2017; September 3 2018; September 4 2018. Published in Journal of Integer Sequences, September 9 2018.

Return to Journal of Integer Sequences home page.